# Canonical decompositions of 3-connected graphs

### v2

@@ 3.3.4. A structural 2-Separation Theorem
If $v\in S(A,B)$, then $(N(v)+v,V(G)-v)$ is a 2-separation of $G$ that crosses $(A,B)$. If $S(A,B)=N(v)$, then $(A,B)$ is crossed by a 2-separation of $G$ which contains $v$ in its separator. ∎
Proof of Theorem 1.6.1.
Let $G$ be a 2-connected graph, and let $N$ denote the set of totally-nested 2-separations of $G$. By Section 3.3.4 and Section 3.3.5, $N$ induces a tree-decomposition ${{\mathcal{T}}}(N)=:(T,{{\mathcal{V}}})$ of $G$. Since $G$ is 2-connected and $(T,{{\mathcal{V}}})$ has adhesion two, it follows with Menger’s theorem that all torsos of $(T,{{\mathcal{V}}})$ are minors of $G$. By Theorem 3.3.2, the torsos of $(T,{{\mathcal{V}}})$ (which coincide with the torsos of $N$) are 3-connected, cycles or $K_{2}$’s.
@@
Let $N^{\prime}$ denote the set of all totally-nested 2-separations of $G^{\prime}$. As $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ cross like in a cycle, they are not members of $N^{\prime}$. By Section 3.3.4, $(A^{\prime},B^{\prime})$ interlaces a unique splitting star $\sigma^{\prime}$ of $N^{\prime}$, and $(C^{\prime},D^{\prime})$ interlaces $\sigma^{\prime}$ as well (since otherwise $(C^{\prime},D^{\prime})$ would be nested with $(A^{\prime},B^{\prime})$). By the (Structural 2-Separation Theorem). (3.3.2), the torso of $\sigma^{\prime}$ is a cycle; let us denote this cycle by $O^{\prime}$. Since $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ cross like in a cycle, $O^{\prime}$ alternates between the separators $S(A^{\prime},B^{\prime})$ and $S(C^{\prime},D^{\prime})$.
Claim 3.3.2.2.
@@
Proof of Claim. Let $(U,W)\in N$. By Section 3.3.2, $(U,W)$ is externally 2-connected. This is preserved by subdivision, so $(U^{\prime},W^{\prime})$ is externally 2-connected, and totally-nested by Section 3.3.2. Hence $(U^{\prime},W^{\prime})\in N^{\prime}$.
Clearly, the map $\varphi$ is injective. It remains to show that it is surjective, so let $(X,Y)\in N^{\prime}$ be given. By Section 3.3.5, the separator of $(X,Y)$ contains no subdividing vertices, so $(U,W)$ where $U:=X\cap V(G)$ and $W:=Y\cap V(G)$ is a 2-separation of $G$ which $\varphi$ sends to $(X,Y)$. As above, applying Section 3.3.2 twice gives $(U,W)\in N$. $\diamondsuit$
By Claim 3.3.2.2, $\sigma:=\varphi^{-1}(\sigma^{\prime})$ is a splitting star of $N$. Since the separators of the elements of $\sigma^{\prime}$ contain no subdividing vertices, the torso $O$ of $\sigma$ is obtained from $O^{\prime}$ be replacing every subpath $xv_{e}y$ where $e=xy\in F\cap E(O^{\prime})$ with the edge $e$. Thus, $O$ is a cycle. As the cycle $O^{\prime}$ alternates between $S(A^{\prime},B^{\prime})$ and $S(C^{\prime},D^{\prime})$, the cycle $O$ alternates between $S(A,B)$ and $S(C,D)$.
@@
We claim that $(A,B)\in N$. If $X$ is 3-connected, then $G[B]$ is 2-connected by Section 3.3.5. If $X$ is a cycle, then $G[A]$ is 2-connected by Section 3.3.5. Hence at least one of $G[A]$ or $G[B]$ is 2-connected. Since $G$ is obtained from $X$ by replacing some edges with connected graphs containing their endvertices, and since the graph $X\setminus A$ is connected, also the graph $G[B\setminus A]$ is connected. Hence $(A,B)\in N$ by Section 3.3.2.
But then $(A,B)$ is an element of $N$ that interlaces $t$ (viewed as a splitting star of $N$), which contradicts Section 2.2.2. ∎

### v3

@@ 4 A structural 2-Separation Theorem
If $v\in S(A,B)$, then $(N(v)+v,V(G)-v)$ is a 2-separation of $G$ that crosses $(A,B)$. ∎
Proof of Theorem 7.1.
Let $G$ be a 2-connected graph, and let $N$ denote the set of totally-nested 2-separations of $G$. By Section 4 and Section 5, $N$ induces a tree-decomposition ${{\mathcal{T}}}(N)=:(T,{{\mathcal{V}}})$ of $G$. Since $G$ is 2-connected and $(T,{{\mathcal{V}}})$ has adhesion two, it follows with Menger’s theorem that all torsos of $(T,{{\mathcal{V}}})$ are minors of $G$. By Theorem 4.1, the torsos of $(T,{{\mathcal{V}}})$ (which coincide with the torsos of $N$) are 3-connected, cycles or $K_{2}$’s.
@@
Let $N^{\prime}$ denote the set of all totally-nested 2-separations of $G^{\prime}$. As $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ cross like in a cycle, they are not members of $N^{\prime}$. By Section 4, $(A^{\prime},B^{\prime})$ interlaces a unique splitting star $\sigma^{\prime}$ of $N^{\prime}$, and $(C^{\prime},D^{\prime})$ interlaces $\sigma^{\prime}$ as well (since otherwise $(C^{\prime},D^{\prime})$ would be nested with $(A^{\prime},B^{\prime})$). By the (Structural 2-Separation Theorem). (4.1), the torso of $\sigma^{\prime}$ is a cycle; let us denote this cycle by $O^{\prime}$. Since $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ cross like in a cycle, $O^{\prime}$ alternates between the separators $S(A^{\prime},B^{\prime})$ and $S(C^{\prime},D^{\prime})$.
Claim 5.0.1.
@@
Proof of Claim. Let $(U,W)\in N$. By Section 2, $(U,W)$ is externally 2-connected. This is preserved by subdivision, so $(U^{\prime},W^{\prime})$ is externally 2-connected, and totally-nested by Section 2. Hence $(U^{\prime},W^{\prime})\in N^{\prime}$.
Clearly, the map $\varphi$ is injective. It remains to show that it is surjective, so let $(X,Y)\in N^{\prime}$ be given. By Section 5, the separator of $(X,Y)$ contains no subdividing vertices, so $(U,W)$ where $U:=X\cap V(G)$ and $W:=Y\cap V(G)$ is a 2-separation of $G$ which $\varphi$ sends to $(X,Y)$. As above, applying Section 2 twice gives $(U,W)\in N$. $\diamondsuit$
By Claim 5.0.1, $\sigma:=\varphi^{-1}(\sigma^{\prime})$ is a splitting star of $N$. Since the separators of the elements of $\sigma^{\prime}$ contain no subdividing vertices, the torso $O$ of $\sigma$ is obtained from $O^{\prime}$ be replacing every subpath $xv_{e}y$ where $e=xy\in F\cap E(O^{\prime})$ with the edge $e$. Thus, $O$ is a cycle. As the cycle $O^{\prime}$ alternates between $S(A^{\prime},B^{\prime})$ and $S(C^{\prime},D^{\prime})$, the cycle $O$ alternates between $S(A,B)$ and $S(C,D)$.
@@
We claim that $(A,B)\in N$. If $X$ is 3-connected, then $G[B]$ is 2-connected by Section 5. If $X$ is a cycle, then $G[A]$ is 2-connected by Section 5. Hence at least one of $G[A]$ or $G[B]$ is 2-connected. Since $G$ is obtained from $X$ by replacing some edges with connected graphs containing their endvertices, and since the graph $X\setminus A$ is connected, also the graph $G[B\setminus A]$ is connected. Hence $(A,B)\in N$ by Section 2.
But then $(A,B)$ is an element of $N$ that interlaces $t$ (viewed as a splitting star of $N$), which contradicts Section 2. ∎
Acknowledgments
We are grateful to Nathan Bowler and Reinhard Diestel for questions that indirectly led us to a streamlining of the definition of tri-separations. We are deeply grateful to the two reviewers for their outstanding support with this overlong paper. We thank the first reviewer for finding errors in the proofs of Section 5, Section 5 and Section 7, and for optimising the proofs of Section 3 and Section 4. We thank the second reviewer for finding a gap in the proof of Section 4. We thank Romain Bourneuf for independently finding the error in the proof of Section 7. Raphael W. Jacobs and Paul Knappe have studied mixed-separations independently from us, and we thank them for interesting discussions on mixed-separations. We thank Dominik Blankenhagen and Nicolás Pich Preuss for spotting typos in the appendix.